40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 244 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  5064   Sat Jul 30 00:33:33 2011 KojiUpdatePSLABSL Laser crystal temp left largely excited & left unattended for more than 3hours

I found that the ref cav trans CCD view was blinking with 30-50 fringe amplitudes. This meant the laser freq was swinging ~50GHz.

I checked the ABSL laser and the SG out of a lock-in amplifier was connected to the slow input.

This was shaking the laser temp from 29degC to 46degC. This was the cause of the fringe swinging.
This big excitation changing the output power too as the temp was changed across it mode-hop region.

I have disconnected the excitation from the laser no matter how useful experiments were took place as there was no e-log entry about this.

I need the explanations

1. Why our precious laser is exposed to such a large swing of temperature?

2. Why the excitation is left like that without any attendance?

3. Why there was no elogging about this activity?

  5063   Fri Jul 29 18:43:02 2011 Manuel, IshwitaUpdatePEMplugging seismometers to ADC

[Manuel, Ishwita, Jenne, Jamie]

We changed the C1PEM model and the names of the C1:PEM channels.

We reinstalled the blue breakout box, since the purple one still didn't work.

So, now the AA board channels are connected as follows...

C1 = C1:PEM-SEIS_GUR1_X

C2 = C1:PEM-SEIS_GUR1_Y

C3 = C1:PEM-SEIS_GUR1_Z

C4 = C1:PEM-SEIS_GUR2_X

C5 = C1:PEM-SEIS_GUR2_Y

C6 = C1:PEM-SEIS_GUR2_Z

C7 = C1:PEM-SEIS_STS_1_X

C8 = C1:PEM-SEIS_STS_1_Y

C9 = C1:PEM-SEIS_STS_1_Z

C11 = C1:PEM-SEIS_STS_2_X

C12 = C1:PEM-SEIS_STS_2_Y

C13 = C1:PEM-SEIS_STS_2_Z

C14 = C1:PEM-SEIS_STS_3_X

C15 = C1:PEM-SEIS_STS_3_Y

C16 = C1:PEM-SEIS_STS_3_Z

C17 = C1:PEM-ACC_MC1_X

C18 = C1:PEM-ACC_MC1_Y

C19 = C1:PEM-ACC_MC1_Z

C20 = C1:PEM-ACC_MC2_X

C21 = C1:PEM-ACC_MC2_Y

C22 = C1:PEM-ACC_MC2_Z

Although the channels for all 3 STS-2 seismometers are made but only one is installed. So only Channels C1 to C9 are now in operation...

We checked the data from the plugged channels with the Dataviewer. We could see the peak whenever someone jumped in the lab. Even Kiwamu jumped and saw his signal.

  5062   Fri Jul 29 16:25:06 2011 kiwamuUpdateASCbeam axis and Y arm aligned

Last night I aligned the incident beam axis and the Yarm by touching the PZT mirrors and the suspensions.

I didn't estimate how good they were aligned, but I guess the Y arm is now ready for the Y green light.

 Next : Y green alignment and the MC spots measurement / alignment.

 

 ++ Motivation ++

Prior to the coming vent we want to have the Y arm, incident beam axis and Y green light aligned so that we can align some necessary optics in the chamber.

Also alignment of the incident beam will allow us to re-position the incident beam alignment monitor (i.e. IPPOS and IPANG).

Our plan was to first align the Y arm using the ASS system and then align the Y green light to the Y arm.

 

++ what I failed ++

First I was trying to measure the spot positions on the MC mirrors to make me sure the beam axis has/hasn't changed.

Also I was going to align the MC suspensions to have nice spot position on each suspension using the MCASS system

because this will help us checking the beam clearance in the Faraday and perhaps re-positioning of the Faraday during the coming vent.

But essentially I failed and eventually gave up because MCASS didn't work. It seems that MCASS needs some modifications in the scripts.

Then, to make me feel better I moved on to the Y arm and beam axis alignment.

 

++ what I did ++

I tried using C1ASS to align the incident beam and suspensions on the Y arm, but it didn't work.

However the drive signals from ASS and its demodulated signals looked fine. Only the feedback did not work correctly.

Every time I enabled the feedback paths, the arm just lost the lock. Something is wrong in the feedback paths.

Then I started to align the cavity by my hands while looking at the demodulated signal from each LOCKIN module.

I aligned the things until each demodulated signal fluctuates around zero.

At the end the beam spots on the ETMY and ITMY camera looked well-aligned and the transmitted light became larger by a factor of 2ish.

  5061   Fri Jul 29 16:04:51 2011 steveUpdateSUSETMX, PRM & SRM OSEM gains set

Jamei fixed the computer. Now I had a healthier ETMX with SIDE signal that allowed me to set gains to get Q of 5

The Y arm restore locked the arm at TRY -out 2.4V,    the X arm still did not lock or align to lock. It has some fringes.

ETMX  POS gain   22  -> 17,              PIT  3  ->  1,                    YAW  3,                     SIDE     -170    ->    -110

PRM                        50  ->  30,                     1,                                       2,                                       50

SRM                        55  ->  20,                    2.4  ->  1.5,                      4.8  ->  2,                          140   ->    100

Now all SUS OSEM gains are set

  5060   Fri Jul 29 12:39:26 2011 jamieUpdateCDSc1iscex mysteriously crashed

c1iscex was behaving very strangely this morning.  Steve earlier reported that he was having trouble pulling up some channels from the c1scx model.  I went to investigate and noticed that indeed some channels were not responding.

While I was in the middle of poking around, c1iscex stopped responding altogether, and became completely unresponsive.  I walked down there and did a hard reset.  Once it rebooted, and I did a burt restore from early this morning, everything appeared to be working again.

The fact that problems were showing up before the machine crashed worries me.  I'll try to investigate more this afternoon.

  5059   Fri Jul 29 12:25:54 2011 Ishwita, ManuelUpdatePEMSTS-2 seismometer box

The 'Bacardi' STS-2 seismometer was tested with the "purple" breakout box and it was found out that all the three axes gave a voltage of 11 V (as shown on the screen of the oscilloscope) before pressing the auto-zero button and after pressing it the voltage shown was 6 V. We tried again the blue box and it was working perfectly after pushing the auto zero button (the auto zero took a few seconds). The power of the purple box is still on, we will wait a few hours, to see if something changes.

  5058   Thu Jul 28 21:52:40 2011 ranaUpdateComputersanother attempt to use pianosa
  1. Pianosa doesn't cache the SVN pwds so you need to re-enter at each SVN up or commit. This is different from the rest of our workstations. We need to determine what behavior we want.
  2. Tried to use the netGPIB scripts:

pianosa:gpib 0> ./readSR785.csh rb2
rb2
netgpibdata.py: Command not found.

  5057   Thu Jul 28 19:49:12 2011 SonaliUpdatePSLPMC trans beam aligned.

Kiwamu and I aligned the PMC transmitted beam the incident beam going to PMC today.

I learnt how to lock the PMC using the digital controls.

  5056   Thu Jul 28 17:04:04 2011 NicoleUpdateSUSPhotosensor Head Calibration Curve

Here is the calibration curve (displacement versus voltage output) for the photosensor head that I made with the S5971 photodiodes and L9337 LEDs. This was made using a regular mirror. The linear region appears to be between 0.4 and 0.75cm. I will need to arrange the photosensor head so it measures displacements in the linear region of this plot. This plot was made using a 287 ohm resistor.

187K_resistor.jpg

  5055   Thu Jul 28 16:26:42 2011 steveUpdateSUSITMX, ITMY & ETMY OSEM gains adjusted

 

OSEM damping gains were adjusted by observing  real time dataviewer to get Q of 5

OSEMs were kicked up one by one with 200 counts  ~1sec. The error signal was optimized to get 1/2 of exitation amplitude at the 5th sinusoid wave.

C1: SUS-ITMX_SUSPOS_N1 gain   111      ->        65,             PIT        7.2      ->    8,           YAW        12      ->     6,          SIDE     280

                  ITMY                                      277       ->     120,                        19.2     ->     7,                             24     ->     19,                        420    ->      470

                  ETMY                                       10        ->       32,                          20      ->     3,                             20     ->     10,                          50

                  ETMX                                       22        ->      25,                             3,                                                3,                                              -170  

ETMX having problems:  1, YAW can not be excited

                                             2, SIDE has no error signal in dataviewer. Sensing voltage on MEDM screen 0.142V

 

                      

 

  5054   Thu Jul 28 16:10:34 2011 kiwamuUpdateLSCBoth arm locked

[Nicole / Jamie / Rana / Kiwamu]

  The X arm and Y arm have been locked.

The settings for the locking were stored on the usual IFO_CONFIGURE scripts, so anybody can lock the arms.

In addition to that Nicole, Jamie and Rana re-centered the beam spot on the ETMY_TRANS camera and the TRY PD.

The next step is to activate the C1ASS servo and align the both arms and beam axis.

 

 

Xarm locking notes:

* Changed TRX gain from -1 to -0.02. Without this 50x reduction the arm power was not normalized.

* Had to fix trigger matrix to use TRX for XARM and TRY for YARM. Before it was crazy and senseless.

* Lots of PZT alignment. It was off by lots.

* Yarm trans beam was clipping on the steering mirrors. Re-aligned. Needs to be iterated again. Be careful when bumping around the ETMY table.

* YARM gain was set to -2 instead of -0.2. Because the gain was too high the alignment didn't work right.

ALWAYS HAVE an OPEN DATAVIEWER with the standard ARM channels going when doing ANY INTERFEROMETER WORK.

THIS IS THE LAW.

  5053   Thu Jul 28 16:00:28 2011 kiwamuUpdateComputer Scripts / Programsanother offset script : offset2

A new offset-zeroing script has been developed and it is ready to run.

 

 The motivation is to replace the old zeroing script called offset by a better one because this old script somehow failed to revert the gain settings on a given filter bank.

The new script, named offset2, does the same job, but uses tdsavg instead of using ezcaservo. So it doesn't screw up the gain settings.

Additionally the structure of the script is much simpler than the old offset script, and fewer ezca-functions.

I will modify some scripts which use the old offset script so that all the offset-zeroing is done by offset2.

 

P.S.

Useful scripts are listed on the 40m wiki

http://blue.ligo-wa.caltech.edu:8000/40m/Computers_and_Scripts/All_Scripts

  5052   Thu Jul 28 13:51:00 2011 SonaliUpdateGreen LockingZHL-32A-S.

Initially I was using RFPD-1611to get the IR beat frequency. Its gain was not very high, so I was getting a very low signal of power -37 dBm.

I used ZHL-32A-S with a gain of 25 dBm to amplify it before feeding it into the spectrum analyser.

I connected the ground of the amplifier circuit to the red of the power supply, which blew the amplifier.

I learnt that there is a small tab indicating the ground side of the BNC to banana connectors which I should have noticed.

I learnt to plug in the side with th little tab on it into the ground of the power supply. (Learnt it the hard way I guess!!)

 

 

  5051   Thu Jul 28 02:33:04 2011 JenneUpdateLockingYarm flashing, but not yet locked

Because I'm too lazy to write a cohenrent elog right now, here's my notes that I wrote while working tonight:

Elog notes, 27July2011

Aligned Xarm, just to check on it.  Had to flip sign of TRX in DCPD filter bank (to gain of -1) to make the signal positive.

Restored Yarm, see some slight flashing, but no lock yet.
Adjusted phase rotation of AS55 from 56.5deg to 60deg, just by-eye trying to maximize AS55I, my arm error signal. AS55I goes from ~ -40 to +60 counts

Tried fitzing with Yarm gain, flipping sign, incr gain. No real change in signals, or flashing.


Incr. ETMY oplev gains to -0.4 from -0.2
Engaged ELP35's on Pit and Yaw, to be more similar to other optics.  However, right now all of the optics have different things in their filter banks.  Why??

Arm is flashing pretty reliably now, but still not locking.  The trigger threshold is always satisfied, so that's not it.

  5050   Wed Jul 27 15:49:56 2011 steveUpdateSAFETYManuel receives safety training

Our surf student Manuel Marchiò received 40m specific safety training today.

Attachment 1: P1080096.JPG
P1080096.JPG
  5049   Wed Jul 27 15:49:13 2011 jamieConfigurationCDSdataviewer now working on pianosa

Not exactly sure what the problem was, but I updated to the head of the SVN and rebuilt and it seems to be working fine now.

  5048   Wed Jul 27 15:42:41 2011 kiwamuSummaryGeneralmitues of 40m meeting : Task list

Quote:

The vent will start from 1 st of August ! 

 ++++ Task List for the vent preparation ++++

  + Preparation of beam dumps (Jamie / Steve)

  + Health check of shadow sensor and measurement of the cross-coupling  (Steve)

  + Measurement of the arm Lengths and estimation of the required precision (Kiwamu)

  + Alignment of the Y green beam (Suresh)

  + Alignment of the incident beam axis (Jenne)

  + Measurement of the MC spot positions  (Suresh)

  + Loss measurement of the arm cavities (Kiwamu / volunteers)

 

 ++++ Task List for the post-vent activity ++++

  + 3f RFPDs (Koji / Rana)

  + EOM resonant circuit (Kiwamu)  

  + Sophistication of the LSC model (Yoichi)

  + DRMI commissioning (Keiko / Anamaria)

 

  5047   Wed Jul 27 15:38:01 2011 steveUpdateVACbeam traps for vacuum

We have cleaned, baked, rga scaned traps for vacuum. Thorlabs LB1 on 1" OD ss posts and forks. The effective surface area is 43 x 18 mm of stacked razors.

Seven pieces are mounted on New Focus #9962  1" OD SS vented-  pedestrals to 5.5" center height and 5 pieces to 4.875"

Attachment 1: P1080095.JPG
P1080095.JPG
  5046   Wed Jul 27 15:18:50 2011 kiwamuSummaryGeneralschedule

The vent will start from 1 st of August ! !

  5045   Wed Jul 27 12:31:47 2011 Manuel, IshwitaSummaryPEMWeekly summary

We kept reading about digital filtering

We tested the seismometer last friday

Jan came and tested again the seismometer last monday

We wrote a simulation of the stacks transfer functions, and of the distance between the mirrors.

 

  5044   Wed Jul 27 12:19:19 2011 NicoleSummarySUSWeekly Summary

Since last week, I've been working on building the photosensor head and have been making adjustments to my photosensor circuit box.

Changes to photosensor circuit (for box):

1) Last week, I was reading in the two signals from the two heads through a single input. Now there are two separate inputs for the two separate photosensors

2)During one of my many voltage regulator replacements, I apparently used a 7915 voltage regulator instead of a 7805 (thanks, Koji, for pointing that out! I never would have caught that mistake X___X)

3)I was powering my 5V voltage regulator with 10V...Now I'm using 15 V (now I only need 1 power supply and 3 voltage input plugs)

I have also began assembling my first photosensor head. Here is what I have so far:

sensorhead.JPG

 

Here is what needs to be done still for the photosensor head

I need to find four Teflon washers and nuts to rigidly attach the isolated PCB (PCB, Teflon sheet combination) to the box. I already have the plastic screws in (I want to use plastic and Teflon for electrical isolation purposes, so as to not short my circuit).

I need to attach the sheath of my signal cable to the box of the photosensor head for noise reduction (plan: drill screw into photosensor head box to wrap sheath wires around)

I need to attach the D-sub to the other end of my signal cable so that it can connect to the circuit box. So far, I only have the D-sub to connect the cable to my photosensor head

Yesterday, Suresh helped to walk me through the photosensor box circuit so that I now understand what voltages to expect for my circuit box trouble-shooting. After this lesson, we figured out that the problem with my photosensor box was that the two op-amps were saturated (so I fixed the feedback!). After replacing the resistor, I got the LED to light up! I still had problems reading the voltage signals from the photodiodes. I was reading 13.5V from the op amp output, but Koji explained to me that this meant that I was too close to saturation (the photodiodes were perhaps producing too much photocurrent, bringing the output close to saturation). I switched the 150 K resistor in the feedback loop to a 3.4K resistor and have thus successfully gotten displacement-dependent voltage outputs (i.e. the voltage output fluctuates as I move my hand closer and farther from the photosensor head). 

Now that I have a successful circuit to power and read outputs from one photosensor, I can begin working on the other half of the circuit to power the other photosensor! 

sensorcircuit.JPG

  5043   Wed Jul 27 10:10:12 2011 steveUpdatePSLlocking status

 80 days: PMC is drifting

Attachment 1: 80dpmcMC.jpg
80dpmcMC.jpg
  5042   Wed Jul 27 10:04:29 2011 SureshUpdateIOOWFS transfer function measurements

This is part of the WFS activity.  So far I have completed the following tasks:

1)  I fixed the MEDM screens up to a point where they can be used for locking.  There are still some buttons which invoke non-existing screens and some blank fields.  But the basic filter banks and input  and output matrices are fixed.

2) I copied all the old filter banks into the new screens both in the WFS head and in the WFS Master, where the servo filters are located.  The I and Q filter banks in the WFS heads have been switched on.

3) I <=> Q phase settings in the WFS head for each quadrant:  We have assumed that the I and Q are orthogonal so D=90 for all cases.  I set the R phase to minimise the signal in all the Q lines.  So the signal is largely in the I phase.  I used Sine Response feature in DTT while supplying an excitation signal to MC2_ASCPIT_EXC.  At times I used the YAW instead of PIT if I did not get enough coherence.  This was set manually by watching the Q phase signal and minimising that by adjusting the R angle.  It was in general possible to get this correct to a deg.   There are several old scripts to do this in the MC/WFS but they do not work since most of them are based on the ezlockin or ezcademod functions.    I will try to fix the ezWFS1phase and ezWFS2phase scripts to automate this.  Some channel names have to be changed in these.

4) I measured the transfer function between the mirror motions [(MC1, MC2, MC3) x (PIT, YAW)] and the sensor DoF [(WFS1, WFS2) x (PIT, YAW)].  The measurements are reported below.  The plan is to invert this matrix and use it as the Out_Matrix.

WFS_TF_Phase_Sheet1.png

I list here the various steps I took in making this measurement.

a) Set the DC offsets on the individual quadrants to zero using an old script (which I updated with the new channel names).  The script is called McWFS_dc_offsets and is located in the $scripts$/MC/WFS directory.   Note that before doing this the PSL shutter was closed.  This script sets a basic EPICS parameter called AOFF for each channel.  These are listed in cvs/cds/caltech/target/c1iool0 .

b) Then the PSL beam into the MC was steered to optimise coupling into MC (described in my earlier post today).  This is because we use the input beam as a reference while setting up the WFS.

c) Unlock the MC and center the directly reflected beam from the MC on the WFS.  We use the DC monitors on the C1IOO_WFS_QPD.adl screen to center the spot on the WFS head. 

d) Then used the WFSoffsets script to set the offsets in the I and Q filter banks to zero.  This script uses the ezcaservo to look at the OUT16 channels and zeroes them by setting an appropriate offset.  I took care to switch off all slow filters in the I and Q filter banks before this operation was carried out .  Only the 60Hz comb filter was on.

e) Opened the PSL shutter and relocked the MC

f) Then I measured the transfer co-efs by oscillating the optic (exciting a specific degree of freedom) and observing the response in the WFS sensor degrees of freedom.   These are tabulated above.

Next

   I plan to use this matrix and prepare the Output matix and then close the WFS servo loops. 

 

 

  5041   Wed Jul 27 08:59:10 2011 SureshUpdateLSClocking status

 I had to realign PSL beam into the MC in order to reobtain the MC lock.  We lost lock at sometime around 8:30 AM on Tuesday.  See attached trend data for MC_RFPD_DCMON. 

The is the second time this week that I had to do this when we were unable to obtain the MC lock.  On both occassions the zig-zag at the end of the PSL table was tweaked to minimise the MC_RFPD_DCMON.

MC_RFPD_DCMon.png

 We have been using the MC as a Beam Axis Reference.  And therefore we are adjusting the PSL beam to maximise coupling into MC.  However if MC's beam axis has shifted, then would it not be best to use the pzt's to re-obtain coupling into the arm cavities? 

Quote from #5040

Beam axis:

 I guess that the beam axis has changed a lot to the horizontal direction.
The beam spots on the REFL and AS camera looked off-centered by a size of the spot.
The beam axis has to be well-aligned before the vent.

  5040   Wed Jul 27 01:58:23 2011 kiwamuUpdateLSClocking status

Through some locking exercise I found that several things are degrading.

Remember the interferometer is like a cat, so we have to feed and take care of her everyday. (Otherwise the cat will be dead !)

 

Beam axis:

 I guess that the beam axis has changed a lot to the horizontal direction.
The beam spots on the REFL and AS camera looked off-centered by a size of the spot.
The beam axis has to be well-aligned before the vent.

 

Locking of the Arms :

 didn't lock at all. It could be a problem of the demodulation phase on AS55.
Also the TRY camera looked pretty much off-centered. The spot is already getting out from the field of view.
We have to fix this issue, otherwise we cannot align the beam axis.
 
 

Locking of PRM :

 Sort of okay, I was able to lock both MICH and PRCL although I had to flip the sign of the MICH control gain due to the demod-phase change.
The suspensions don't look healthy. The beam spots on the REFL and AS camera move a lot even without any length feedback.
It means some of the suspensions are shaky.
  5039   Wed Jul 27 01:57:28 2011 SonaliUpdateGreen LockingWeekly Summary

1. I have used the PMC  trans beam in my set-up as the required PSL beam.

2. I have superposed the ETMX-Fibre output with the PSL beam on the PSL table.

3. I have used suitable beam splitters and lens to match the power and the  sizes of the overlapping beams and have aligned them to the optimum.

4. A lens having f=7.6 cms is used to focus the beam into the PD.

5. Initially, I used the broadband 1611 NewFocus PD to find the IR beat signal by scanning the oven temperature. (using the digital sitemap controls.)

6. I checked the previous elog entries by Suresh and Koji on the green beat signal they had worked on and used their data to get an idea of the temperature range of the oven where I could obtain a beat.

7. I obtained peaks at three different temperatures as had been noted previously and set the temperature so that I am now sitting in the middle stable regime.

8. Then I switched to the 1811 100 MHz PD as it has a larger gain. It has a saturation power of 100 microWatts. The input power at the PD is measured to be 80 microWatts.

9. I was having trouble getting a clean peak due to presence of many harmonics as seen on the spectrum analyser. This happened because there was too much power incident on the PD which led to arising of non-linearity giving rise to harmonics.

10.To reduce the power entering the PD, I put in a ND 1.0 Filter just before the beam enters the PD and obtained a clean signal.

11. I will use  the frequency counter tomorrow to check the resonant frequency and try to connect the output to acquire a digital signal.

12. Otherwise I will proceed to build a Mixer Frequency Discriminator.

13. After the feed-back loop is completed, I will proceed to compare the frequency-noises of the green-beat lock and the IR-beat lock.

  5038   Tue Jul 26 21:11:40 2011 haixingSummaryElectronicsAA filter tolerance analysis

Given this new setup, we realized that the previous tolerance analysis is incorrect. Because the uncertainty in the capacitance value
does not affect the common mode rejection, as two paths share the same capacitor. Now only the imbalance of two resistors is relevant.
The error propagation formula goes as follows:


We require that the common-mode rejection error at low frequency up to 8kHz, namely
with , one can easily find out that the corner frequency needs to be around 24kHz.


 

  5037   Tue Jul 26 11:38:45 2011 steveUpdateSUSETMY bias checked

 

 I test drove ETMY biases.

PITCH  worked well in slow and fast modes. Slow drive was from the IFO alignment screen C1:SUS-ETMY_PIT_COM and

the fast one from C1:SUS-ETMY_ASCPIT_OFFSET

YAW did not.  It was always diagonal. It was  specially bad with the fast drive. I compared them with ETMX. ETMX yaw is diagonal a little bit too.

The OPLEV return spots on the qpd ETMX and ETMY are big 5-6 mm diameter. The ETMY spot has weird geometry to qpd.

  5036   Tue Jul 26 09:01:53 2011 JennyUpdateComputer Scripts / ProgramsMode matching

I found a mode matching solution to match the beam coming to the PSL table from the AP table so that I can lock the laser beam coming onto the PSL table to the reference cavity on the table. I determined that at the polarizing beam splitter, I want a beam with a q=(147+25.1i)mm (w0=58mm). This came from applying the ABCD matrices for three distances,

  • d1=693 mm,
  • d12=660.4 mm, and
  • d2=393.7 mm, separated
  • an f=229.1 mm planoconvex lens and
  • an R=300 mm curved mirror.

to a beam with q0 = 406.4i mm (w0=0.371 mm at the PMC).

I obtained the following mode matching solution, which I will try to implement on the PSL table:

The beam I have has waist 0.281 mm at -2.74 m (I set my origin at the polarizing beam splitter--the spot where I want my beam to match the beam coming from the PMC, so all waists are behind that point). These numbers  come from the beam-profiling and MATLAB-fitting I did (see 5015).

The solution I chose was: f = 1145.6 mm at -0.95 m and f = 572.7 mm at -0.62 m. This may need to be changed however, if I need to add in some beam steering, which would increase the path length traveled by the beam.

modematchparameters.png modematchpic.png

 

  5035   Tue Jul 26 03:15:52 2011 JenneUpdateLockingLatest MICH noise budget

[Jenne, Rana]

We had another look at the MICH noise budget tonight. Rana has verified that my techniques / math aren't too ridiculous. 

In the first attachment, you'll notice that the MICH noise is waay above the shot noise of 1mW on the beam splitter.  We don't know why.  One problem is that the modulation depth of the 55MHz is too low by ~a factor of 10.  Kiwamu and his magical resonant circuit are working on fixing this.  This will not, however, fix the huge discrepancy here.  More investigation and meditation is required!  For this measurement, the whitening gain of AS55 was set to 42dB for both I and Q.

In the 2nd attachment, the PSL shutter is closed, so all of these are dark measurements of AS55.  (The input matrix on the LSC screen is AS55Q * 1 -> MICH_IN1, so they're the same).  All we've done is change the whitening gain before the ADC.  For 0dB and 9dB, you can see that the low freq noise didn't change - here we're still limited by the ADC noise.  With 21dB and 42dB we're clear of the ADC, so either is fine.  Unfortunately, the high freq stuff when the loop is on matches up with the high freq part of the dark noise, so that's part of the problem....

Attachment 1: MICHnoise_shotNoise_25July2011.pdf
MICHnoise_shotNoise_25July2011.pdf
Attachment 2: MICH_darkNoise_whiteningGainChanging_25July2011.pdf
MICH_darkNoise_whiteningGainChanging_25July2011.pdf
  5034   Mon Jul 25 23:43:20 2011 ManuelHowToElectronicsManual for 1201 Low Noise Preamplifier

I found the manual for the Low Noise Preamplifier Model 1201 at this link and I attached it.

The one we have in the lab (S/N 48332) miss the battery packs and miss also the remote programming options input/output. Its inside battery compartment is empty and I found 2 unscrewed screws with washers and nuts inside the preamplifier box. The battery cable are disconnected and they have 2 green tape labels (-) and 2 red tape label (+).

 

 

Attachment 1: ITHACO_1201_Instruction_&_Maintenance.pdf
ITHACO_1201_Instruction_&_Maintenance.pdf ITHACO_1201_Instruction_&_Maintenance.pdf ITHACO_1201_Instruction_&_Maintenance.pdf ITHACO_1201_Instruction_&_Maintenance.pdf ITHACO_1201_Instruction_&_Maintenance.pdf ITHACO_1201_Instruction_&_Maintenance.pdf ITHACO_1201_Instruction_&_Maintenance.pdf ITHACO_1201_Instruction_&_Maintenance.pdf
Attachment 2: DSC_3249.png
DSC_3249.png
  5033   Mon Jul 25 18:51:38 2011 ManuelUpdatePEMSTS-2 seismometer hardware testing with Jan

[Jan, Manuel, Jenne]

Jenne called Jan to check and figure out why the Streckeisen seismometer (SN #100151) doesn't work, hence we checked the output of the seismometer boxes as we did last friday. (This is the problem of seeing the X and Y channels saturated, when we look at them on a floating 'scope, as in the linked elog entry.)

Jan unplugged and plugged again the orange cable into the seismometer and nothing happened.   Well, what Jan was listening for was "clicks" inside the seismometer indicating that it was receiving power.  We heard these, and moved on to examining the breakout boxes.  Also, we checked that we could hear the "clicks" (one per mass) when we pushed the mass-centering button on the little green companion box.

We weren't sure that the purple box was working properly, so since we had seen the blue box work last time, we changed the purple box with the blue box in rack 1X6.

The Z-channel of the purple box returns a correct signal, that means that all the masses in the seismometer work (because the Z-signal is a linear combination of the three masses U, V, W), the X and Y channel have a DC component of about 10 Volts, Jan said that the recentering of the seismometer masses could need all the night, so we keep the power of the box on. If tomorrow morning the X and Y signal won't  both be zero mean, we will open and check the box.

The power of the box is still on so that the masses can recenter overnight.

Edits by JD

  5032   Mon Jul 25 17:16:02 2011 JamieUpdateSUSNow acquiring SUSXXX_IN1_DQ channels

> And.....we have also lost the DAQ channels that used to be associated with the _IN1 of the SUSPOS/PIT/YAW filter modules. Please put them back; our templates don't work without them.

I have (re?)added the SUS{POS,PIT,YAW,SIDE}_IN1_DQ channels.  I did this by modifying the activateDQ.py script to always turn them on [0].  They should now always be activated after the activateDQ script is run.

[0] This script now lives in the cds_user_apps repo at cds/c1/scripts/activateDQ.py

 

  5031   Mon Jul 25 13:09:39 2011 JamieUpdateCDSc1ioo Make problem

> It looks like something wrong with Makefile.

Sorry, this was my bad.  I was making a patch to the makefile to submit back upstream and I forgot to revert my changes.  I've reverted them now, so everything should be back to normal.

  5030   Mon Jul 25 13:01:24 2011 kiwamuUpdateCDSc1ioo Make problem

[Suresh / Kiwamu]

HELP US Jamieeeeeeee !! We are unable to compile c1ioo.

 

It looks like something wrong with Makefile.

We ran make c1ioo -- this was successful every time. However make install-c1ioo doesn't run.

The below is the error messages we got.

        make install-target-c1ioo
        make[1]: Entering directory `/opt/rtcds/caltech/c1/core/branches/branch-2.1'
        Please make c1ioo first

Then we looked at Makefile and tried to find what was wrong. Then found the sentence (in 36th line from the top) saying

        if test $(site)no = no; then echo Please make $$system first; exit 1; fi;\

We thought the lack of the site-name specification caused the error.

So then we tried the compile it again with the site name specified by typing

     export site=c1

in the terminal window.

It went ahead a little bit further, but it still doesn't run all through the Make commands.

 

  5029   Mon Jul 25 11:46:28 2011 steveUpdateSUSETMY sus problem hunt

[Kiwamu / Steve]

We checked some electronics noise on the ETMY shadow sensor system.

Noise from the WF, AA board and ADC are below the shadow sensor spectra on ETMY.

It means something funny is going on in the upstream side (including the satelight box and shadow sensors)

OR the coil drivers side are going crazy ??

 

As Rana pointed out in his entry (#5025), the spectra of the shadow sensors on ETMY were quite bad below 3 Hz. The floor are higher than that of ETMX by factor of 10 or so.

To check if the noise comes from some of the electronics, we disconnected D15-sub from pd to whitening in.

The spectra with/without shadow sensors are attached below.

The curves in brown and green are the ones taken when the shadow sensors were disconnected from the WF board.

So these two curves represent the summed noise of the WF, AA and ADC.

This tells us to look toward the OSEM.

 

Attachment 1: ETMY_AA_ADC.png
ETMY_AA_ADC.png
  5028   Mon Jul 25 11:06:38 2011 kiwamuSummaryGeneralWeekly report

Summary of the week ending July 24th.  Number of elog entries = 45


- LSC
 * Check of LSC WF switching
  -> some were switching, but the majority were not.

- SUS
 * Ran activateDQ.py for seeting some DQ channels of Oplevs
 * Turned ON all the offset buttons on the OL1, etc.
 * Rebuilt and restarted c1msc, c1sus, c1scx and c1scy as an update.
 * ETMY's shadow sensors look bad. Unknown noise below 3 Hz, which is higher than the usual floor by factor of 10.

- ABSL
 * The frequency lock was down.
 * The laser power into the RFPD had been too big, so it was reduced

- OAF
 * Seismometers were connected to the AA-board on 1X6
 * Most of the channels were acquired to the ADC, but some were not.

- Mode Cleaner
 * Gain of quadrants were checked.
 * Due to the SUS model update, the MC locking trigger hasn't worked correctly. This was fixed by changing ioo.db file

- Misc.
 * Virtual box was installed on Rossa. Altium is now available on Rossa.

  5027   Mon Jul 25 11:04:22 2011 kiwamuSummaryGeneralWeekly report

Summary of the week ending July 17th.  Number of elog entries = 20

- LSC
 * BO switching logic for the WF was installed on c1lsc
 * Channel mapping updated

- SUS
 * Oplev health check. Spectrum of each quadrant on every suspension was inspected and looked healthy.

- OAF
 * BNS interface board was attached to an AA board
 * The AA board was installed on 1X7. The Electro-optic fanout chassis on 1X7 is now sitting on a jack, this should be fixed.

- Fiber experiment (ALS)
 * Fibre from the ETMX and EMTY tables were routed to the PSL table

- Misc.
 * Alberto came over to the 40m with Wagonga
 
  5026   Mon Jul 25 11:02:19 2011 kiwamuSummaryGeneralWeekly report

 Summary of the week ending July 10th.  Number of elog entries = 21

- SUS

 + The cutoff frequency of the high pass filters for the damping were set to 30Hz.
 + Turned off all the BounceRoll filters.
 + The BS oplev was checked and seemed healthy.
 

- LSC


 + All the measred data of the LSC whitening filters were fit.
 + All the zpk parameters are recorded on the wiki.

- ABSL

 

+ The setup completed.  
 + The freqeucy-lock of the ABSL laser was achieved with UGF of ~ 40kHz.
 + The temperature of the ABSL laser was adjusted to be 47.25 deg
 

- ALS

 (Fiber experiment)
 + The I-P curve of the ETMY laser was measred.
 + The current set point is 1.8 [A], which used to be 1.5 [A], corresponding to the output of power of 197 [mW] and 390 [mW] respectively.
 

  5025   Mon Jul 25 00:35:44 2011 ranaUpdateSUSsomething wrong with ETMY LR sensor

a.png

Looks like either the LR OSEM is totally mis adjusted in its holder or the whitening eletronics are broken.

Also looks like the ETMY is just not damped at 1 Hz? How can this be?

I look at the SUS_SUMMARY screen which apparently only Steve and I look at:

bad.png

Looks like the suspensions have factor of 10-100 different gains. Why?

**  The ETMY just doesn't behave correctly when I bias it. Both pitch and yaw seem to make it do yaw. I leave this for Jamie to debug in the morning.

***  Also, the BIAS buttons are still broken - the HOPR/LOPR limits ought to be 5000 and the default slider increment be 100. Also the YAW button readback doesn't correctly show the state of the BIAS.

****  And.....we have also lost the DAQ channels that used to be associated with the _IN1 of the SUSPOS/PIT/YAW filter modules. Please put them back; our templates don't work without them.

  5024   Sun Jul 24 22:19:19 2011 haixingSummaryElectronicsAA filter tolerance analysis

 

>> This sort of OK, except the capacitor connects across the (+) terminals of the two input opamps, and does not connect to ground:

low_pass_new.png

 

>> Also, we don't care about the CMRR at 64 kHz. We care about it at up to 10 kHz, but not above.

In this case, the conner frequency for the low-pass filter would be around 100kHz in order to satisfy the requirement.


>>And doesn't the value depend on the resistors?

Yes, it does. The error in the resistor (typically 0.1%)  is much smaller than that of the capacitor (10%). Since the resistor error propagates in the same as the capacitor,
we can ignore it.

Note that we only specify the conner frequency (=1/RC) instead of R and C specifically from the tolerance analysis, we still need to choose appropriate
values for R and C with the conner frequency fixed to be around 100kHz, for which we need to consider the output impedance of port 1 and port 2.

 

 

  5023   Sun Jul 24 20:47:21 2011 ranaSummaryElectronicsAA filter tolerance analysis

This is sort of OK, except the capacitor connects across the (+) terminals of the two input opamps, and does not connect to ground.

Also, we don't care about the CMRR at 64 kHz. We care about it at up to 10 kHz, but not above. The sample frequency of the ADC is 64 kHz, but all of the models run at 16 kHz or less, so the Nyquist frequency is 8 kHz.

And doesn't the value depend on the resistors?

  5022   Sun Jul 24 20:36:03 2011 haixingSummaryElectronicsAA filter tolerance analysis

Koji and Haixing,

We did a tolerance analysis to specify the conner frequency for passive low-pass filtering in the AA filter of Cymac. The
link to the wiki page for the AA filter goes as follows (one can have a look at the simple schematics):
http://blue.ligo-wa.caltech.edu:8000/40m/Electronics/BNC_Whitening_AA

Basically, we want to add the following passive low-pass filter (boxed) before connecting to the instrumentation amplifier:

low_pass.png

Suppose (i) we have 10% error in the capacitor value and (ii) we want to have common-mode rejection
error to be smaller than 0.1% at low frequencies (up to the sampling frequency 64kHz), what would be
conner frequency, or equivalently the values for the capacitor and resistor, for the low-pass filter?

Given the transfer function for this low-pass filter:
transfer_function.png     f0.png
and the error propagation equation for its magnitude:
error.png
we found that the conner frequency needs to be around 640kHz in order to have
DT.pngwith Dc.png
 



  5021   Sat Jul 23 02:24:10 2011 SureshUpdateIOOrestarted the frame builder

I restarted the fb twice during the last 15mins.   This was after I added test points into the C1IOO/WFS1.mdl and C1IOO/WFS2.mdl.

  5020   Fri Jul 22 17:01:41 2011 Iron ManFrogsGeneralProof that Alberto lived through his Iron Man!

Alberto_IronMan_small.jpgIronman Vineman 70.3 logo

print 208 Alberto Stochino 67/129 585/1376 36:02 1:52 830 6:41 2:38:58 21.1 296 4:58 56:33 2:13:40 - 5:40:19
 
  5019   Fri Jul 22 15:39:55 2011 haixingUpdateSUSmatching the magnets

Yi Xie and Haixing,

We used the Gauss meter to measure the strength distribution of bought magnets, which follows a nice Gaussian distribution.
We pick out four pairs--four fixed magnets and four for the levitated plate that are matched in strength. The force difference is
anticipated to be within 0.2%, and we are going to measure the force as a function of distance to further confirm this.

In the coming week, we will measure various transfer functions in the path from the sensors to the coils (the actuator). The obtained
parameters will be put into our model to determine the control scheme. The model is currently written in mathematica which can
analyze the stability from open-loop transfer function.

  5018   Fri Jul 22 14:22:13 2011 Ishwita, ManuelUpdatePEMSTS-2 seismometer hardware testing

We have two STS-2 seismometer boxes... the blue box & the purple box. Initially we used the blue box for the STS-2 seismometer (named Bacardi by Jenne).

  • Oscilloscope powered on battery was used to test the blue box by observing the velocity output of the three axes (X, Y, Z).  It was found out that the mean value of DC volt of...

X = +10 V

Y = +11 V

Z = -0.1 V

Thus, X and Y axes showed abnormally high DC volt. It was also found out that in AC coupling mode of the oscilloscope... changes were observed in the signal received from Z axis when some seismic wave was generated near the Bacardi by jumping near it. No such changes were observed from signals received from X & Y axes.

  • We removed the blue box and used the purple box for the same Bacardi seismometer & used the oscilloscope powered on battery to test it. It was found out that the mean value of DC volt of...

X = +4.4 V

Y = +4.4 V

Z = +4.4 V

In Ac coupling mode of the oscilloscope... changes were observed in the signals from X, Y, Z axes when someone jumped near Bacardi.

  • The above voltages from the two STS-2 seismometer boxes are unsuitable for the ADC box since it works with voltages ranging from +2 V to -2 V.... meaning it will consider any voltage signal above +2 V as +2 V and any signal below -2 V as -2 V. Hence we need to find out how to use these STS-2seismometer boxes with the ADC box.
  • We also tried measuring the DC volt from the shield and the center of a BNC connector corresponding to Y axis of the purple box (lets call it 'BNC-test') by using BNC-to-banana adaptors and banana wires. Signal from shield of BNC-test was sent to oscilloscope's channel 1 (connected to center of its BNC connector) and signal from center of BNC-test was sent to oscilloscope's channel 2 (connected to center of its BNC connector). On the oscilloscope screen it was observed that both the signals gave the same mean voltage output (-2.2V).
  5017   Fri Jul 22 10:24:34 2011 steveUpdateVACpressure plot at day 213

Dec 21, 2010 we pumped down the MARK4 rebuilt 40m-IFO and the malev has been pumping on it since than

Attachment 1: p.jpg
p.jpg
  5016   Fri Jul 22 10:13:41 2011 steveUpdateVACRGA scan at day 213
Attachment 1: pd70m213scan.jpg
pd70m213scan.jpg
  5015   Thu Jul 21 23:36:51 2011 JennyUpdate Fitting beam waist with MATLAB

I am starting work on the PSL table at the 40m. My goal is to lock the laser coming from the nearby table to the FP cavity and get a measurement of the response to a temperature step on the surrounding can.

I have to mode match the beam to the cavity. Specifically, I have to mode match to the beam coming from the PMC through the EOM to the polarizing beam splitter. Yesterday David and I measured the beam width at various distances (from a particular lens through which the beam traveled), and I fit that data using MATLAB to find the beam's waist size and location. However, I'm not convinced that the fit is any good, since we only took measurements at five spots and they had large error bars.

 

z (mm) 2w_vert (mm) 2w_horiz (mm)
180 4.68 3.38
230 4.64 3.49
305 4.68 3.47
370 5.1 3.81
510 5.5 4.17

Here is the fit I obtained using fminsearch. The horizontal beam width measurements were smaller than the vertical width measurements, suggesting that the incoming beam was elliptical. I fit the data for each set of measurements separately and got two waist locations. The red trace is the fit for the horizontal width and the blue represents the vertical width of the beam. Averaging the two fitted waist locations and sizes gives

vert z_0= -1760 mm (waist location)

horiz z_0= -1540 mm (waist location)

vert w_0 = 0.286 mm (waist size)

horiz  w_0 = 0.275 mm (waist size)

avg z_0= -1650 mm

avg w_0 = 0.281 mm

 

twobeamfit2.jpg

Here is the code I used:

I defined the function spotsize.m and then made a function gaussbeam.m that called it with input parameters and returned the least squares error. I then wrote another function twobeamfits.m that ran fminsearch to minimize the least squares error and made the above plot. I've pasted the code below.

spotsize

function omega = spotsize(z_0, w_0, z)
lambda=0.001064;
omega=w_0*(1+(lambda*(z-z_0)/(pi*w_0^2)).^2).^(1/2);

 

gaussbeam

function sse = gaussbeam(params,xvals,yvals)

%This f'n takes as its inputs
%three parameters (w_0, z_0, and lambda),
%a vector of x-values (distances),
%and an associated vector of y-values (spotsizes),


%It then generates a vector of fitted y-values by applying
%an exponential approach function (single pole), with the given parameters,
%to the x-values.

%It then returns the sum of the squares of the entries of the difference
%between the fitted y-vector and the actual y-vector

z_0=params(1);
w_0=params(2);
fityvals=spotsize(z_0, w_0, xvals);

error=(fityvals - yvals);% .*xvals;
% sse stands for sum of squares error
sse=sum(error.^2);

 

twobeamfits

function [outputs] = twobeamfits(guesses, dists, vert, horiz)


%This f'n takes as its inputs
%two starting guess parameters (w_0 and z_0),
%a vector of distances (x-values),
%and two associated vectors of measured beam radii,

%the radius measured along the vertical axis

%and the radius measured along a horizontal axis (y-values).

%It then calls the gaussbeam f'n for each set of y-values and minimizes its output (sum of squares error)
%using the fminsearch f'n. It outputs the fit parameters it settles on.

%It then plots the input data, the fitted curves, and the residuals


fminopts=optimset('TolFun',1e-6,'MaxIter', 100000);
vertparams=fminsearch(@gaussbeam,guesses,fminopts,dists,vert);
fitvert=spotsize(vertparams(1), vertparams(2), dists);
resid1=(vert-fitvert)./vert;
spoterror=[.1, .1, .1, .1, .1]; %uncertainties, all in mm

fminopts=optimset('TolFun',1e-6,'MaxIter', 100000);
horizparams=fminsearch(@gaussbeam,guesses,fminopts,dists,horiz);
fithoriz=spotsize(horizparams(1), horizparams(2), dists);
resid2=(horiz-fithoriz)./horiz;


points=linspace(-2000,1000,1000);
figure(1)
hold off
clf
subplot(2,1,1)
hold on
errorbar(dists, vert, spoterror, 'x')
grid
errorbar(dists, horiz, spoterror, 'r*');
plot(points,spotsize(vertparams(1), vertparams(2), points));
plot(points,spotsize(horizparams(1), horizparams(2), points),'r');
xlabel('Distance z (mm)')
title('Gaussian Beam Fits')
ylabel('Spotsize w (mm)')
legend('Vertical Spotsize','Horizontal Spotsize','Vertical Fit',...
    'Horizontal Fit','Location','SouthEast')
hold off

subplot(2,1,2)
plot(dists,resid1,'x')
hold on
plot(dists,resid2,'r*');
xlabel('Distance (z)')
title('Residuals')
ylabel('Fractional Difference')
legend('Vertical Fit Residuals','Horizontal Fit Residuals',...
    'Location','SouthEast')
grid

outputs=[vertparams horizparams];

 

 

 

Later on I may repeat some measurements and try to gain more certainty in my fit. In the mean time I will use this beam profile for mode matching. 

 

ELOG V3.1.3-